OpenStudy (anonymous):

Compute the monthly payments for an add-on interest loan of $850, with an annual interest rate of 15 percent and a term of 4 years. Round to the nearest cent as needed.

OpenStudy (chaise):

Similar to your last question. A=PiT A=850*0.15*4 A= 510 The interest on the loan is 510, plus the original amount of 850. 850+510=1360. There are 48 months in a year. 1360/48=$28.34cents The monthly repayment is $28.34 cents.

OpenStudy (kropot72):

"add-on interest loan" implies compound interest to me rather than simple interest.

OpenStudy (chaise):

I see. haha, in his last question I took the assumption that it was simple interest. :\

OpenStudy (kropot72):

Assuming monthly compounding of the interest the amount to be paid over four years is: \[850\times(1+(.15/12))^{48}=$1,543.05\] Monthly payments = $1,543.05/48 = $32.15